Argon has three naturally occurring isotopes, argon-36 (atomic mass=35.97 u), argon-38 (atomic mass=37.96 u), and argon-40 (atomic mass= 39.96 u).The natural abundance of argon-36 is 0.334%, and abundance of argon-38 is 0.063%.                                                                   Determine the average atomic mass of argon.  Show all work. 

1
Expert's answer
2021-05-07T05:28:38-0400

Abundance of Argon-40=100-(0.334+0.063)

=99.603%

0.334%=0.00334

0.063%=0.00063

99.603%=0.99603

Average Atomic Mass=[(35.97*0.00334)+(37.96*0.00063)+(99.603*39.96)]

=39.8014+0.1201+0.0239=39.9454amu

=39.95amu